# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v2

@@ 6 Conclusion
Acknowledgments
We would like to thank Bruce Lee, Celestine Mendler-Dünner, and Lars Lorch for feedback on early versions of the paper. We also thank Reese Pathak and Pierre Bizeul for helpful discussions. JH was supported by the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. PW was supported by the Max Planck ETH Center for Learning Systems. AS was supported by the Swiss National Science Foundation under grant 204439. GK conducted the initial part of this work during his visit to the IDEAL Institute, hosted by Lev Reyzin, which was supported by NSF ECCS-2217023.
@@ Acknowledgments
appendix.Asubsection.A.1subsection.A.2appendix.Bassumption.1assumption.2assumption.3assumption.5appendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dappendix.Esubsection.E.1subsection.E.2subsection.E.2subsection.E.3subsection.E.3equation.E.23subsection.E.4subsection.E.5subsection.E.6subsection.E.6subsection.E.6subsection.E.6subsection.E.6subsection.E.6table.caption.19subsection.E.7
@@ A.1 Limitations of global learning under the LRH
We consider the Gaussian measure $\gamma_{d}$ on $\mathbb{R}^{d}$ and define the family of $s$-sparse, $k$-locally linear functions (with $k$ “cells”), denoted by $\mathcal{F}_{d,s,k}\subset\{\mathbb{R}^{d}\to\mathbb{R}\}$, as defined as
$\mathcal{F}_{d,s,k}=\bigg\{f^{\star}(x)=\sum_{i=1}^{k}w_{i}^{\top}x\cdot\mathbf{1}_{K_{i}}:\ \forall$ $1\leq i<j\leq k\;\gamma_{d}(K_{i}\cap K_{j})=0,\|w_{i}\|_{0}\leq s,\|w_{i}\|_{2}\lesssim 1$ (4)
$\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1,\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim\frac{s}{d}\bigg\},.$
@@
Proposition (informal, see Proposition 10 below).
Assume that $\sigma\lesssim 1$, $k\asymp\mathrm{Poly}(d)$, and $s\asymp 1$}. Then Equation 5 holds. Furthermore, for $n\gtrsim\exp(\Omega((\log d)^{2}))$, there exists a polynomial-time algorithm $\mathcal{A}$ such that
@@
We refer to the survey of Reyzin (2020) for more details on the low-degree polynomial conjecture. We emphasize that our assumptions on $\mathcal{F}_{d,s,k}$ are much less restrictive than imposing a uniform bounded Lipschitz (or an average $L_{2}$ Lipschitz constraint), as the Lipschitz constant may be high on the boundary between two cells, and the results of Bizeul & Klartag (2025) implying that Lipschitz functions can be learned with polynomial number of samples.
Roughly speaking, from a geometric view, this spectral concentration implies that most of the energy of the coefficients is localized at the decision boundaries between the cells $\{K_{i}\}$. An “underparameterized” global classifier, in this case, the best low-degree approximation, necessarily smooths these boundaries. Therefore, by isoperimetry in high dimensions, this smoothing leads to significant overlap between the decision boundaries of cells. Since most samples fall in the areas of these “distorted/smoothed” decision boundaries, we need many samples to learn such functions. [^8] It is well-known that even for $1$-Lipschitz function, its best low-degree polynomial (it terms of $L_{2}$), is highly non-Lipschitz cf. Bizeul & Klartag (2025). The latter aligns with previous observations regarding the required complexity (and the lack of robustness) of foundation models (e.g., Bubeck & Sellke, 2021).
@@ Appendix B Formal assumptions
Next, we quantify the expressivity of the learned features $\Psi$. Naturally, features needs to be sufficiently expressive for any linear model in feature space to approximate the ground truth function. First, we make the relatively straightforward assumption that locally, the learned features are a linear recombination of concepts.
@@
While standard TTT omits the explicit constraint, we use this formulation to analyze the specialization mechanism we observe empirically. It remains to state standard assumptions:
@@
Finally, we assume that $\mathbf{\Psi}_{x^{\star}}$ satisfies the generalized restricted eigenvalue (GRE) condition. Combined with local linearity (Assumption˜4), the GRE is simply the standard restricted eigenvalue condition on the local concept design matrix $\mathbf{\Phi}_{x^{\star}}$, a condition that is fundamental for guaranteeing stable recovery in sparse regression (Bickel et al., 2009; Van De Geer & Bühlmann, 2009).
@@
$\mathbb{E}[(y^{\star}-\langle\Psi(x^{\star}),\hat{v}_{x^{\star}}^{\text{TTT}}\rangle)^{2}]\leq\sigma^{2}+O\left(\frac{\sigma^{2}s\log(d_{1}/s)}{k}\right)+O(s\,\eta_{\text{mis}})+\eta_{\text{inherent}}.$
@@ C.1 Proof of Proposition 2
$\mathbb{E}[(y^{\star}-\langle\Psi(x^{\star}),\hat{v}_{x^{\star}}^{\text{TTT}}\rangle)^{2}]\leq\sigma^{2}+\left(\mathcal{E}_{\text{inherent}}+\mathcal{E}_{\text{estimation}}\right)^{2}$
@@
Let $h=\tilde{v}_{x^{\star}}-\hat{v}_{x^{\star}}^{\text{TTT}}$. Since by Assumption˜5, both $\tilde{v}_{x^{\star}}$ and $\hat{v}_{x^{\star}}^{\text{TTT}}$ satisfy the constraint of Equation˜6, the error vector $h$ is sparse in the concept space: $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$.
@@
Step 3: We analyze the right hand side of Equation˜9. Using local linearity (cf. Assumption˜4), $h^{\top}\mathbf{\Psi}_{x^{\star}}^{\top}=(P_{x^{\star}}^{\top}h)^{\top}\mathbf{\Phi}_{x^{\star}}^{\top}$. Let $Z=\frac{1}{k}\mathbf{\Phi}_{x^{\star}}^{\top}(\Delta+\varepsilon)$ be the total concept score. Then,
@@
Substituting this back into Equation˜9 gives:
@@
Step 4: Applying GRE. Since $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$, we can apply the GRE (cf. Assumption˜9) to Equation˜10:
@@
Combining this with Equation˜11 gives:
@@
Substituting the definitions from Lemmas 5 and 6:
@@ C.2 Concentration bounds
Hence by Dudley’s entropy integral (applied with the $L_{2}$ metric on $\mathcal{U}_{m}$ and $\operatorname{diam}(\mathcal{U}_{m})\leq 2$),
@@ C.4 Hardness of global learning under the LRH: main proposition and proof
Let $c_{1}\geq 0$, $\sigma\lesssim 1$, $k\asymp d^{c_{1}}$, and $s\asymp 1$. Assume $n\gtrsim\exp(\Omega(\log d)^{2})$. Then, Equation 5 holds, and there exists a polynomial-time algorithm $\mathcal{A}$ in its input, such that
@@
Using Equation 22, we conclude that for $m\in 1,\ldots,c\log(d)$ (by choosing $t_{m}\asymp 1/m$)
@@
which cannot align with the previous equation and our assumption of $\|f^{\star}\|_{L_{2}(\gamma_{n})}=1$. In words, this property says that $f^{\star}$ is far from being a “low degree” polynomial. Meaning that,
@@
Therefore, we need to learn at most the top $\Theta(\log(d))$ basis of the Hermite polynomials. Using standard estimates from approximation theory (cf. Kalai et al. (2008)), samples to learn the $m=\Theta(\log(d))$-degree polynomials, let’s say by the projection estimator, which in words means to learn the coefficients of the best low degree polynomial Bizeul & Klartag (2025).
Hence, we need at most
$n\gtrsim s^{2}\log(k)\cdot d^{C\log(k)}\asymp s^{2}\log(k)\cdot k^{\log(d)s^{2}}\asymp\exp(C\log(d)^{2})$
samples to learn $f^{\star}\in\mathcal{F}_{d,s,k}$, and the claim follows.
By definition, one can easily see that $\mathcal{F}_{d,s,k}$ is much harder to learn than the Gaussian Index Model (GIM). As the generative component of our functions are $k^{*}\asymp\log(d/n)$. It implies that under the low-polynomial degree conjecture, the aforementioned classes require at least $\exp(\Theta(\log(d/n)^{2}))$ samples to even learn GIM, see Damian et al. (2024) and references within, and the seminal work of Arous et al. (2021).
@@
$\{1_{K}(x)\quad\text{$K$ is polytope in $\mathbb{R}^{d}$ with $\mathrm{Poly}(d)$ facets}\}\subset\{\mathbb{R}^{d}\to\mathbb{R}\}$
requires at least $\exp(\Omega(\log(d))\asymp\mathrm{Poly}(d)$ samples in the information theoretic sense. However, the algorithmic gap exists in those classical works Klivans et al. (2008); Kalai et al. (2008), and a gap in the matching upper bound for the sample complexity.
Appendix D Additional ablations
While the improvement of TTT in image classification may seem small in terms of classification error, we find that TTT can significantly reduce cross-entropy loss on the test set, as shown in Figure˜6.
@@ Appendix D Additional ablations
Appendix E Experiment details
@@
E.1 Comparison of logits (Figure˜3)
We start by introducing the related notation. Define the top-$10$ probabilities for TTT in the sparse concept space for the test point $x^{\star}$ as a vector $p^{\hat{\Phi}}(x^{\star})\in[0,1]^{10}$, similarly, $p^{\hat{\Psi}}(x^{\star})$ for dense reconstructions. In this notation, we implicitly assume that the probabilities $p^{\hat{\Phi}}(x^{\star})$ are scaled by some optimal factor $\tau_{x^{\star}}$. In particular, define the corresponding to $p^{\hat{\Phi}}(x^{\star})$ logits (i.e., before the softmax application) by $\mathrm{logit}^{\hat{\Phi}}(x^{\star})$. Then for each test point $x^{\star}$ we “calibrate” (cf. (Xie et al., 2024)) the concept space probabilities $p^{\hat{\Phi}}(x^{\star})$ by finding the optimal scale $\tau_{x^{\star}}$ which aligns them closer. Namely, we choose $\tau_{x^{\star}}$ to minimize the following KL-divergence term:
@@ E.1 Comparison of logits (Figure˜3)
$\bar{p}^{\hat{\Phi}}(x^{\star})=\mathrm{softmax}\left(\frac{\mathrm{logit}^{\hat{\Phi}}(x^{\star})}{\tau_{x^{\star}}}\right).$
We note that although the temperature $\tau_{x^{\star}}$ is tweaked on a sample-by-sample basis, the variation is not significant. In particular, we observe the mean temperature of $\approx 0.8$ (predictive distribution $p^{\hat{\Psi}}(x^{\star})$ in dense space has a bit “sharper” profile as $\tau_{x^{\star}}<1$) and standard deviation of $\approx 0.1$.
In these terms, the relative total variation at position $i\in\left\{1,\dots,10\right\}$ is defined as follows:
@@
E.2 SAE on MNIST data
This setup is evaluated on MNIST (1010k test samples, 6060k training samples). We use a Gemma-Scope-style SAE (Lieberum et al., 2024) to obtain a sparse concept vector $\smash{\hat{\Phi}(x_{i})}$ for each image $x_{i}\in\mathcal{D}_{\mathrm{MNIST}}$. The only conceptual difference to regular applications of SAEs is that our encoder has a convolutional architecture, i.e., is LeNet-like (LeCun et al., 1998a) and maps each image to a representation vector $\Psi(x_{i})\in\mathbb{R}^{d_{2}}$ with $d_{2}=256$, which is then lifted to a sparse concept vector $\smash{\hat{\Phi}(x_{i})}$. However, the decoder is still linear to force linearity of the concept space. We also directly reconstruct the inputs, i.e., images $x$, which means that reconstruction error takes the following form:
$\mathbb{E}_{x\in\mathcal{D}}\|x-D\cdot\hat{\Phi}(x)\|_{2},\quad D\in\mathbb{R}^{28\cdot 28\times d_{1}}$
In this setup, $\smash{\hat{\Phi}(x_{i})\in\mathbb{R}^{d_{1}}}$ where $d_{1}=1024$ and the average sparsity of the concept vector

### v3

@@ 6 Conclusion
appendix.Asubsection.A.1subsection.A.2appendix.Bassumption.1assumption.2assumption.3assumption.5appendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dsubsection.D.1subsection.D.2appendix.Eappendix.Fsubsection.F.1subsection.F.2subsection.F.2subsection.F.3subsection.F.3equation.22subsection.F.4subsection.F.5subsection.F.6subsection.F.6subsection.F.6subsection.F.6subsection.F.6subsection.F.6table.caption.22subsection.F.7
@@ A.1 Limitations of global learning under the LRH
We consider the Gaussian measure $\gamma_{d}$ on $\mathbb{R}^{d}$ and introduce the family of $s$-sparse, $k$-locally linear functions (with $k$ “cells”), denoted by $\mathcal{F}_{d,s,k}\subset\{\mathbb{R}^{d}\to\mathbb{R}\}$, defined as
$\mathcal{F}_{d,s,k}=\bigg\{f^{\star}(x)=\sum_{i=1}^{k}w_{i}^{\top}x\cdot\mathbf{1}_{K_{i}}$ $:\ \forall 1\leq i<j\leq k\;\gamma_{d}(K_{i}\cap K_{j})=0,\|w_{i}\|_{0}\leq s,$ (3)
$\|w_{i}\|_{2}\lesssim 1,\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1,\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim\frac{s}{d}\bigg\}.$
@@
Proposition 2 (informal, see Proposition 11 below).
Assume that $\sigma\lesssim 1$, $k\asymp\mathrm{Poly}(d)$, and $s\asymp 1$}. Then Equation˜4 holds. Furthermore, for $n\gtrsim\exp(\Omega((\log d)^{2}))$, there exists a polynomial-time algorithm $\mathcal{A}$ such that
@@
We refer to the survey of Reyzin (2020) for more details on the low-degree polynomial conjecture. We emphasize that our assumptions on $\mathcal{F}_{d,s,k}$ are much less restrictive than imposing a uniform bounded Lipschitz constraint (or an average $L_{2}$ Lipschitz constraint), as the Lipschitz constant may be high on the boundary between two cells, and the results of Bizeul & Klartag (2025) show that Lipschitz functions can be learned with a polynomial number of samples.
Roughly speaking, from a geometric view, this spectral concentration implies that most of the energy of the coefficients is localized at the decision boundaries between the cells $\{K_{i}\}$. An “underparameterized” global classifier, in this case, the best low-degree approximation, necessarily smooths these boundaries. Therefore, by isoperimetry in high dimensions, this smoothing leads to significant overlap between the decision boundaries of cells. Since most samples fall in the areas of these “distorted/smoothed” decision boundaries, we need many samples to learn such functions. [^8] It is well-known that even for $1$-Lipschitz function, its best low-degree polynomial (it terms of $L_{2}$), is highly non-Lipschitz, cf. Bizeul & Klartag (2025). The latter aligns with previous observations regarding the required complexity (and the lack of robustness) of foundation models (e.g., Bubeck & Sellke, 2021).
@@ Appendix B Formal assumptions
Next, we quantify the expressivity of the learned features $\Psi$. Naturally, features need to be sufficiently expressive for any linear model in feature space to approximate the ground truth function. First, we make the relatively straightforward assumption that locally, the learned features are a linear recombination of concepts.
@@
While standard TTT omits the explicit constraint, we use this formulation to analyze the specialization mechanism that we observe empirically. It remains to state standard assumptions:
@@
Finally, we assume that $\mathbf{\Psi}_{x^{\star}}$ satisfies the generalized restricted eigenvalue (GRE) condition. Combined with local linearity (Assumption˜4), the GRE is simply the standard restricted eigenvalue condition on the local concept design matrix $\mathbf{\Phi}_{x^{\star}}$, a condition that is fundamental to guarantee stable recovery in sparse regression (Bickel et al., 2009; Van De Geer & Bühlmann, 2009).
@@
$\mathbb{E}\left[\big(y^{\star}-\langle\Psi(x^{\star}),\hat{v}_{x^{\star}}^{\text{TTT}}\rangle\big)^{2}\right]\leq\sigma^{2}+O\left(\frac{\sigma^{2}s\log(d_{1}/s)}{k}\right)+O(s\,\eta_{\text{mis}})+\eta_{\text{inherent}}.$
@@ C.1 Proof of Proposition 3
$\mathbb{E}\left[\big(y^{\star}-\langle\Psi(x^{\star}),\hat{v}_{x^{\star}}^{\text{TTT}}\rangle\big)^{2}\right]\leq\sigma^{2}+\left(\mathcal{E}_{\text{inherent}}+\mathcal{E}_{\text{estimation}}\right)^{2}$
@@
Let $h=\tilde{v}_{x^{\star}}-\hat{v}_{x^{\star}}^{\text{TTT}}$. Since by Assumption˜5, both $\tilde{v}_{x^{\star}}$ and $\hat{v}_{x^{\star}}^{\text{TTT}}$ satisfy the constraint of Equation˜5, the error vector $h$ is sparse in the concept space: $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$.
@@
Step 3: We analyze the right hand side of Equation˜8. Using local linearity (cf. Assumption˜4), $h^{\top}\mathbf{\Psi}_{x^{\star}}^{\top}=(P_{x^{\star}}^{\top}h)^{\top}\mathbf{\Phi}_{x^{\star}}^{\top}$. Let $Z=\frac{1}{k}\mathbf{\Phi}_{x^{\star}}^{\top}(\Delta+\varepsilon)$ be the total concept score. Then,
@@
Substituting this back into Equation˜8 gives:
@@
Step 4: Applying GRE. Since $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$, we can apply the GRE (cf. Assumption˜9) to Equation˜9:
@@
Combining this with Equation˜10 gives:
@@
Substituting the definitions from Lemmas 6 and 7:
@@ C.2 Concentration bounds
Hence, by Dudley’s entropy integral (applied with the $L_{2}$ metric on $\mathcal{U}_{m}$ and $\operatorname{diam}(\mathcal{U}_{m})\leq 2$),
@@ C.4 Hardness of global learning under the LRH: main proposition and proof
Let $c_{1}\geq 0$, $\sigma\lesssim 1$, $k\asymp d^{c_{1}}$, and $s\asymp 1$. Assume $n\gtrsim\exp(\Omega(\log d)^{2})$. Then Equation˜4 holds and there exists a polynomial-time algorithm $\mathcal{A}$ in its input, such that
@@
Using Equation˜21, we conclude that for $m\in 1,\ldots,c\log(d)$ (by choosing $t_{m}\asymp 1/m$)
@@
which cannot align with the previous equation and our assumption of $\|f^{\star}\|_{L_{2}(\gamma_{n})}=1$. In words, this property says that $f^{\star}$ is far from being a “low degree” polynomial. Meaning that
@@
Therefore, to obtain a $0.1$ approximation to $f^{\star}$, we need to learn the coefficients of at most (and at least) the top $\Theta(\log(d))$ basis of the Hermite polynomials. Using the result of Bizeul & Klartag (2025) (or the classical work of Kalai et al. (2008)), it can be done with
$n\lesssim s^{2}\log(k)\cdot d^{C\log(k)}\asymp s^{2}\log(k)\cdot k^{\log(d)s^{2}}\asymp\exp(C\log(d)^{2})$
samples. By definition, one can easily see that $\mathcal{F}_{d,s,k}$ is much harder to learn than the Gaussian Index Model (GIM). Since the generative component of our functions satisfies $k^{*}\asymp\log(d/n)$, it follows that, under the low-degree polynomial conjecture, these classes require at least $\exp(\Theta(\log(d/n)^{2}))$ samples to even learn the GIM, see Damian et al. (2024) and references within, and the seminal work of Arous et al. (2021).
@@
$\{1_{K}(x):\text{$K$ is polytope in $\mathbb{R}^{d}$ with $\mathrm{Poly}(d)$ facets}\}\subset\{\mathbb{R}^{d}\to\mathbb{R}\}$
requires at least $\exp(\Omega(\log(d))\asymp\mathrm{Poly}(d)$ samples in the information theoretic sense. However, an algorithmic gap remains in these classical works Klivans et al. (2008); Kalai et al. (2008), and there is also a gap in the corresponding upper bound for the sample complexity.
Appendix D Additional experimental results
This section provides additional experimental results that complement the findings presented in the main body.
D.1 Scaling with feature dimension
Figure 6: Scaling model width. Classification error rates when scaling model width. We evaluate a globally trained model (black) across different model sizes, as well as TTT (red) and a majority vote on the neighborhood (gray). While majority vote leads to a poor predictor with many classes (i.e., complex tasks), TTT consistently outperforms global training, with the performance gap shrinking as the model size increases.
For the scaling experiments on MNIST and ImageNet shown in Figure˜4, we varied only the network width while keeping the depth fixed, and we reported the resulting error rates as a function of the total number of model parameters. Our theoretical analysis (§5) relies on the expressivity of the feature space, which requires the feature dimension $d_{2}$ to scale at least proportionally to $s\log d_{1}$. To make this connection explicit, we present the same results from the main body as a function of the feature dimension in Figure˜6 (instead of the total parameter count). The resulting curves show similar qualitative and quantitative behavior: TTT consistently outperforms global training and majority vote, with the performance gap narrowing as the hidden dimension increases.
D.2 End-to-end test-time training
Because full end-to-end training on the neighborhood set is computationally expensive, it is common to fine-tune only the last linear layer (or the last few layers) as a trade-off between computational cost and performance. An alternative is to apply a parameter-efficient fine-tuning method such as LoRA (Hu et al., 2022). In the experiments in Figure˜4 for the image tasks (MNIST, ImageNet), we adopted the former approach and fine-tuned only the classification head. For the language modeling experiments, we used LoRA to fine-tune the model. Both strategies introduce a capacity constraint on the TTT model. To test whether our conclusions still hold when removing this constraint, we conduct an end-to-end TTT experiment on MNIST.
Figure 7: Classification error rates when scaling model width on MNIST. We compare last-layer TTT (red) against end-to-end TTT (blue) with the globally trained model (black) serving as the baseline. Both TTT variants yield nearly identical error rates.
Mirroring the scaling experiment in Figure˜4, we train LeNet (LeCun et al., 1998a) convolutional neural networks (CNNs) of varying sizes across five random seeds. Details about the global pre-training procedure are provided in Section˜F.4. Based on these pre-trained models, we compare two variants of TTT. First, we apply last-layer TTT, where all layers except the final one are frozen and only the classification head is fine-tuned on the neighborhood set. This matches the setup used in Figure˜4. Second, we perform end-to-end TTT: neighbors are selected using the $L_{2}$-distance in the last-hidden-layer representations of the pre-trained model (as before), but subsequently all model weights are fine-tuned. The resulting model is then used to predict the class label of the test sample. The fine-tuning hyperparameters for end-to-end TTT are provided in Table˜3, obtained by tuning on the validation set.
Hidden dimension
25 50 80 128 176 240 320 512 832 1152 1792
Learning rate 5e-3 0.01 0.01 0.01 5e-3 1e-3 1e-3 1e-3 1e-3 5e-4 5e-4
Epochs 200 200 300 100 200 200 200 300 300 300 200
Neighbors 300 50 200 300 300 50 50 300 300 50 300
Table 3: Hyperparameters for end-to-end TTT.
Appendix E Additional ablations
While the improvement of TTT in image classification may seem small in terms of classification error, we find that TTT can significantly reduce cross-entropy loss on the test set, as shown in Figure˜8.
@@ Appendix E Additional ablations
Appendix F Experiment details
@@
F.1 Comparison of logits (Figure˜3)
We start by introducing the related notation. Define the top-$10$ probabilities for TTT in the sparse concept space for the test point $x^{\star}$ as a vector $p^{\hat{\Phi}}(x^{\star})\in[0,1]^{10}$, similarly, $p^{\hat{\Psi}}(x^{\star})$ for dense reconstructions. In this notation, we implicitly assume that the probabilities $p^{\hat{\Phi}}(x^{\star})$ are scaled by some optimal factor $\tau_{x^{\star}}$. In particular, define the logits corresponding to $p^{\hat{\Phi}}(x^{\star})$ (i.e., before the softmax application) by $\mathrm{logit}^{\hat{\Phi}}(x^{\star})$. Then, for each test point $x^{\star}$, we “calibrate” (cf. (Xie et al., 2024)) the concept space probabilities $p^{\hat{\Phi}}(x^{\star})$ by finding the optimal scale $\tau_{x^{\star}}$ which aligns them closer. Namely, we choose $\tau_{x^{\star}}$ to minimize the following KL-divergence term:
@@ F.1 Comparison of logits (Figure˜3)
$\bar{p}^{\hat{\Phi}}(x^{\star}):=\mathrm{softmax}\left(\frac{\mathrm{logit}^{\hat{\Phi}}(x^{\star})}{\tau_{x^{\star}}}\right).$
We note that although the temperature $\tau_{x^{\star}}$ is adjusted on a sample-by-sample basis, the variation is not significant. Specifically, we find the temperature has a mean of around $0.8$ and a standard deviation of around $0.1$. This mean value ($\tau_{x^{\star}}<1$) indicates that the predictive distribution $p^{\hat{\Psi}}(x^{\star})$ in the dense space has a slightly sharper profile than the baseline.
We now introduce a metric to quantify the remaining discrepancy between the calibrated sparse distribution and the dense distribution. The relative total variation at position $i\in\left\{1,\dots,10\right\}$ is defined as follows:
@@
F.2 SAE on MNIST data
We evaluate the SAE setup on the MNIST dataset (1010k test samples, 6060k training samples). To obtain a sparse concept vector $\smash{\hat{\Phi}(x_{i})}$ for each image $x_{i}\in\mathcal{D}_{\mathrm{MNIST}}$, we employ a Gemma-Scope-style SAE (Lieberum et al., 2024). The only conceptual difference from regular SAE applications is that our encoder has a convolutional architecture, i.e., it is LeNet-like (LeCun et al., 1998a) and maps each image to a representation vector $\Psi(x_{i})\in\mathbb{R}^{d_{2}}$ with $d_{2}=256$, which is then lifted to a sparse concept vector $\smash{\hat{\Phi}(x_{i})}$. However, the decoder is still linear to force linearity of the concept space. We also directly reconstruct the inputs, i.e., images $x$, which means that reconstruction error takes the following form:
$\mathbb{E}_{x\in\mathcal{D}}\|x-D\cdot\hat{\Phi}(x)\|_{2},\quad D\in\mathbb{R}^{28\cdot 28\times d_{1}}.$
In this setup, $\smash{\hat{\Phi}(x_{i})\in\mathbb{R}^{d_{1}}}$ where $d_{1}=1024$ and the average sparsity of the concept vector is